Test whether the series ∞Σn=0 1/(n^5+x^3) converges uniformly or not

1
Expert's answer
2022-03-09T15:26:52-0500

Solution


For the given series


"\sum_{0}^{\infty }\frac{1}{n^5+x^3}"


We use the Ratio Test for the Radius of convergence


Convergence when "L < 1"


"L=\lim_{n\rightarrow \infty }\left | \frac{a_{n+1}}{a_{n}} \right |"


Here, "a_{n}=\frac{1}{n^5+x^3}" , then


"a_{n+1}=\frac{1}{(n+1)^5+x^3}"


Therefore,

"L=\lim_{n\rightarrow \infty }\left | \frac{\frac{1}{(n+1)^5+x^3}}{\frac{1}{n^5+x^3}}\right |"


"L=\lim_{n\rightarrow \infty }\left | \frac{n^5+x^3}{(1+n)^5+x^3}\right |"



"L = \mathop {\lim }\limits_{n \to \infty } \left| {\frac{{{n^5}\left( {1 + \frac{{{x^3}}}{{{n^5}}}} \right)}}{{{n^5}{{\left( {\frac{1}{n} + 1} \right)}^5} + {x^3}}}} \right|\"


"L = \mathop {\lim }\limits_{n \to \infty } \left| {\frac{{{n^5}\left( {1 + \frac{{{x^3}}}{{{n^5}}}} \right)}}{{{n^5}\left( {{{\left( {\frac{1}{n} + 1} \right)}^5} + \frac{{{x^3}}}{{{n^5}}}} \right)}}} \right|\"


"L = \left| {\frac{{\left( {1 + \frac{{{x^3}}}{{{\infty ^5}}}} \right)}}{{\left( {{{\left( {\frac{1}{\infty } + 1} \right)}^5} + \frac{{{x^3}}}{{{\infty ^5}}}} \right)}}} \right|\"


"L = \left| {\frac{{\left( {1 + 0} \right)}}{{\left( {{{\left( {0 + 1} \right)}^5} + 0} \right)}}} \right|\"


"L=1"


Hence the series may be divergent, conditionally convergent, or absolutely convergent.


Now, when


"\sum\limits_0^\infty {\left| {{a_n}} \right|} \" converges, then "\sum\limits_0^\infty {{a_n}} \" converges.


Therefore, the series is convergent for "x>-1"